# Data-driven Discovery for Robust Optimization of Semiconductor Nanowire Lasers

### v1

@@ Abstract
Supporting information is available for this paper, which includes the growth details for each type of nanowire. This can be found at DOI: ????.
This work was funded by UKRI under grants MR/T021519/1, EP/V036343/1 and
EP/W002302/1, by the Deutsche Forschungsgemeinschaft within the frame of the collaborative research center CRC 1375 (project C5), by the Australian Research Council, the Australian National Fabrication Facility ACT node, as well as China special Grant (134000-E62201ZJ); Zhejiang Provincial Natural Science Foundation of China (Z24F040009), Zhejiang University Education Foundation Qizhen Scholar Foundation (K20240015), and CMOS Special Program of Zhejiang University (04010000-K2A033208). Research data supporting this publication will be made available at DOI: (DOI to be provide in due course), and the code to perform the analysis will be made available at DOI: (DOI to be provide in due course), and on github: (URL to be provide in due course).
CRediT author statement: Stephen Church: Formal analysis, Investigation, Methodology, Software, Visualisation, Writing - original draft. Francesco Vitale, Aswani Gopakumar, Nikita Gagrani,Yunyan Zhang, Nian Jiang, Hark Hoe Tan, Chennupati Jagadish, Huiyun Liu, Hannah Joyce and Carsten Ronning: Resources, Writing - review and editing. Patrick Parkinson: Conceptualization, Data curation, Funding acquisition, Methodology, Software, Supervision, Writing - review and editing.

### v2

@@ Abstract
Supporting Information
Supporting information is available for this paper, which includes the growth details for each type of nanowire, some further experimental details, low power PL spectra and statistics of the NW diameters and lifetimes. This can be found at DOI: ????.
Research data supporting this publication is available at DOI: 10.48420/26403496, and the code to perform the analysis will be made available at DOI: (DOI to be provide in due course), and on github: (URL to be provide in due course).
Acknowledgements
This work was funded by UKRI under grants MR/T021519/1, EP/V036343/1 and EP/W002302/1, by the Deutsche Forschungsgemeinschaft within the frame of the collaborative research center CRC 1375 (project C5), by the Australian Research Council, the Australian National Fabrication Facility ACT node, as well as China special Grant (134000-E62201ZJ); Zhejiang Provincial Natural Science Foundation of China (Z24F040009), Zhejiang University Education Foundation Qizhen Scholar Foundation (K20240015), and CMOS Special Program of Zhejiang University (04010000-K2A033208).
CRediT author statement: Stephen Church: Data curation, Formal analysis, Investigation, Methodology, Software, Visualisation, Writing - original draft. Francesco Vitale, Aswani Gopakumar, Nikita Gagrani,Yunyan Zhang, Nian Jiang, Hark Hoe Tan, Chennupati Jagadish, Huiyun Liu, Hannah Joyce and Carsten Ronning: Resources, Writing - review and editing. Patrick Parkinson: Conceptualization, Data curation, Funding acquisition, Methodology, Software, Supervision, Writing - review and editing.